Kerfuffle Corporation is considering the purchase of a new computer system. The cost for the new system, net of set-up and delivery costs, will be $1.6 million. The new system will provide annual before-tax cost savings of $500,000 for the next five years. The increased efficiency of the new system will lower net working capital by $200,000 today. The CCA rate on the new system will be 30%. At the end of five years, the system can be salvaged for $100,000. The firm’s required rate of return is 15%, and its marginal tax rate is 35%. What is the NPV of this cost-cutting project?

Acceptable answers are:

a. -$224,011.86
b. -$22,882.55
c. $15,174.10
d. $76,552.80
e. $563,744.59

** Another expert answered this same question with excellent formatting and detail; however, the answer does not match any of the acceptable choices, so I am attempting to figure out why this is. There is also another "answer" available, but it is illegible and does not actually answer the question. Thanks! **
Answer #1

The New System will provide $500000 before tax cost savings for next five years. Thus the

PV of Cost Saving =PV[(1-Tc)(R-C)] = (1-.35)*$500000/.15(1-(1/1.15)5 ) =$1089450

The New Working Capital Decreases by $200000. Thus

PV[ΔNWC] = -$200000+$200000/(1.15)5 = -$100565

PV of net Capital Spending

PV[NCS] = 1600000 -100000/(1.15)5   =1550282

PV of CCA Tax Sheild = TcdA(1+0.5k)/(k+d)(1+k) - Tc dS /(k+d)(1+k)n

                                      = [.35*.30*1600000*1.075/.45*1.15] -[.35*.30*100000/.45*(1.15)5]

                                      = 348986-11601 = 337385

Thus NPV = PV[(1-Tc)(R-C)]- PV[ΔNWC]- PV[NCS]+PVCCATS

                 = $1089450--$100565 -1550282+337385

                 =-22882

As we have round off the decimal to nearest 1 dollar the answer is Option b